OpenStudy (iesmee):

The surface area of a sphere with radius 7 inches is A. 616 square inches B. 616 cubic inches C. 154 square inches D. 1437 cubic inches

OpenStudy (iesmee):

@skullpatrol

OpenStudy (iesmee):

@dpasingh

OpenStudy (iesmee):

@GoldPhenoix

OpenStudy (goldphenoix):

Do you know the formula for finding the surface area of a sphere?

OpenStudy (iesmee):

no

OpenStudy (anonymous):

Surface Area of a Sphere = 4 pi r^ 2 =\[=4 \times \frac{22}{7} \times 7^2 =4 \times 22 \times 7=616 sq inches \]

OpenStudy (iesmee):

oh that makes sense i have about 6 more of these soo...thanks

OpenStudy (goldphenoix):

The formula for finding the surface area of a sphere is: \[\large \large 4 \pi r^2\]